Calculate how much money a prospective homeowner would need for closing costs on a house that costs $83,500. Calculate based on a 19% down​ payment, 1.5 discount points on the​ loan, a 0.7 point origination​ fee, and ​$1,550 in other fees.

The Closing costs would be:
Answer #1

House price (P) = 83,500

Loan amount = P*(1-downpayment rate) = 83,500*(1-19%) = 67,635

Closing cost = (discount points + origination fee)*Loan amount + other fees

= (1.5% + 0.7%)*67,635 + 1,550 = 3,037.97